OpenStudy (anonymous):

Express the given quantity as a single logarithm. 1/5 ln(x + 2)^5 + 1/2 [ln x − ln(x^2 + 3x + 2)^2]

OpenStudy (ghazi):

\[\frac{ 1 }{ 5 }\log(x+2)^5+\frac{ 1 }{ 2 }[{\ln x- \ln (x^2+3x+2)^2}]= \frac{ 1 }{ 5 }*5 \log(x+2)+\frac{ 1 }{2 }[\ln x- \ln(x^2+3x+2)^2]\] i am sorry to say, this can't be unified because ln and log have different bases

OpenStudy (anonymous):

All of them are base e, surely?

OpenStudy (ghazi):

oops my bad

OpenStudy (anonymous):

Are there any nice identities for\[\log(a)^b\]? Maybe I'm being stupid.

OpenStudy (anonymous):

Also- does anyone know how to have latex flow with your text rather than indent?

OpenStudy (ghazi):

no you are right ...and it can be simplified

jhonyy9 (jhonyy9):

do you know the property of logarithm ? so like ln a + ln b = ? or ln a -ln b = ? or ln a^2 = ?

jhonyy9 (jhonyy9):

you dont like cooperating nothing ? i like help you here ... come on !!!

OpenStudy (anonymous):

i'm confused

jhonyy9 (jhonyy9):

why ? so you know that ln a + ln b = ln a*b yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than do you know how many will be ln a^2 =

jhonyy9 (jhonyy9):

or ln a - ln b =

OpenStudy (anonymous):

ln(a/b)

jhonyy9 (jhonyy9):

yes and ln a^2 =

OpenStudy (anonymous):

2ln a

jhonyy9 (jhonyy9):

yes right sure so than these property of logarithms can us in case of your exercise too ?

OpenStudy (anonymous):

(ln(x + 2) (1/2 ln x / 2ln(x^2 + 3x + 2))

OpenStudy (anonymous):

[ln(x + 2)] [1/2( ln x / 2ln(x^2 + 3x + 2))]

jhonyy9 (jhonyy9):

so the first part is right but the secondly you need to separate 1/5 ln(x+2)^5 = 5/5 ln(x+2) = ln(x+2) this is right so in the second part there are 1/2 (ln x - ln(x^2 +3x +2)^2) = 1/2 (ln (x/(x^2 +3x +2)^2) = ln (x/(x^2 +3x +2)^2)^(1/2) do you understand till now ? can you continue it ?

OpenStudy (anonymous):

yess

jhonyy9 (jhonyy9):

so than how will be ?

OpenStudy (anonymous):

[ln(x+2)] [ln (x/(x^2 +3x +2)^2)^(1/2)]

jhonyy9 (jhonyy9):

why ? than you know that (a^2)^(1/2) =a^(2/2) =a so than what will be ?

jhonyy9 (jhonyy9):

and you know that ln a +ln b = ln a*b

OpenStudy (anonymous):

[ln(x+2)] [ln (x/(x^2 +3x +2)]

jhonyy9 (jhonyy9):

not is right please check it because there is just (x^2 +3x +2)^2 and x on numerator have not exponent 2 so what sign that not can be simplified by 1/2 OK ?

jhonyy9 (jhonyy9):

do you understand it sure ?

OpenStudy (anonymous):

no

jhonyy9 (jhonyy9):

so the first part is right sure will be ln(x+2) OK . in the second part there are 1/2 (ln x - ln(x^2 +3x +2)^2) = 1/2 (ln (x/(x^2 +3x +2)^2)) = ln (x^(1/2) /((x^2 +3x +2)^2)^1/2 = ln (x^1/2) /(x^2 +3x +2) so can you continue it now ?

jhonyy9 (jhonyy9):

so than will result there ln (x+2) + ln (x^1/2) /(x^2 +3x +2) so from this can you continue it ?

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

so than will be using the property of logarithm log a +log b = log a*b so than will be ln ((x+2)*x^1/2 ) /(x^2 +3x +2)

jhonyy9 (jhonyy9):

so x^2 +3x +2 how can you factoriz it ?

jhonyy9 (jhonyy9):

do you know ?

jhonyy9 (jhonyy9):

so x^2 +3x +2 = (x+2)(x+1) is right ?

jhonyy9 (jhonyy9):

so than will be ln (x+2)x^(1/2) /(x+2)(x+1) so simplifie by (x+2) and will result ln x^1/2 /(x+1) so do you understand it now sure ?